OpenStudy (anonymous):

A triangle has an area of 20 cm2. If its base is 40% of its height, what is its height to the nearest tenth of a centimeter?

OpenStudy (anonymous):

say the base= b height = h b=40% h = 0.4h area of triangle = half of the base x height 10 = ........ complete the rest

OpenStudy (anonymous):

Area of a triangle = (1/2)(base)(height) 20 = (1/2)(.4h)(h) 20 = .2h^2 h^2 = 100 h = 10

OpenStudy (anonymous):

Thanks to both of you.

OpenStudy (anonymous):

welcome.